# Robust Amortized Bayesian Inference with Self-Consistency Losses on Unlabeled Data

### v4

@@ 4.1 Multivariate normal model
We first illustrate the usefulness of our proposed self-consistency loss on a controllable toy problem [38]. The prior and likelihood are given by
[…]
The parameters $\theta\in\mathbb{R}^{D}$ are sampled from a $D$-dimensional multivariate normal distribution with mean vector $\mu_{\rm prior}$ and diagonal covariance matrix $\sigma^{2}_{\rm prior}\,I_{D}$. Here, we fix $\mu_{\rm prior}=0$ and $\sigma^{2}_{\rm prior}=1$. On this basis, $K$ independent, synthetic data points $x^{(k)}\in\mathbb{R}^{D}$ are sampled from a $D$-dimensional multivariate normal distribution with mean vector $\theta$ and diagonal covariance matrix $\sigma^{2}_{\rm lik}\,I_{D}$. We fix $\sigma^{2}_{\rm lik}=K$ such that the total information in $x$ remains constant, independent of $K$, which simplifies comparisons across observations of varying number of data points. More details on the training setup and employed neural architectures can be found in Appendix B.
@@ Results
In Figure 1, we depict the results obtained from (a) standard NPE (trained on the simulation-based loss only), (b) our semi-supervised NPE (with the self-consistency loss on known likelihood), and (c) the gold-standard (analytic) reference. We see that standard NPE already completely fails for $x_{\rm obs}\sim N(\mu_{\rm obs}=2,0.01I_{D})$, and subsequently also for any larger values $\mu_{\rm obs}>2$. In contrast, adding the self-consistency loss to obtain our semi-supervised approach achieves almost perfect posterior estimation. This holds true even in cases where $x_{\rm obs}$ is multiple standard deviations away from all the training data, that is, from both the labeled dataset $\{(\theta_{n},x_{n})\}_{n=1}^{N}$ and the unlabeled dataset $\{{x^{*}_{m}}\}_{m=1}^{M}$. These results indicate that the self-consistency criterion can provide strong robustness gains even far outside the typical space of training data.
@@
Figure 2: Posterior distance quantified by maximum mean discrepancy (MMD) to the analytic posterior for variations of the default configuration. Errorbars show ±$\pm 1$ SDs over 10 model refits.
In Figure 2, we report the maximum mean discrepancy (MMD) between the approximate and true posterior the factors parameter dimensionality and number of unlabeled observations. When varying the parameter dimensionality (Figure 2(a)), including the self-consistency loss yields nearly perfect posterior approximation up to 10 dimensions, even with extreme deviations from the initial training data. It also significantly improves accuracy in the 100 parameter scenario. The dataset size factor (Figure 2(b)) shows robust gains, with clear improvements over the standard simulation-based loss even when using as few as four unlabeled observations (versus 1024 labeled ones). In Figure 6 (Appendix C), we additionally report posterior mean and standard deviation bias as well as maximum mean discrepancy for all the above factors. Varying the mean $\mu_{\rm obs}$ of the new observations shows that, as long as the data used for evaluating the self-consistency loss is not identical to the training data (i.e., as long as $\mu_{\rm obs}\neq 0$), including the self-consistency loss component enables accurate posterior approximation far outside the typical space of the training data.
In Figure 7 in Appendix C, we see that the benefits of self-consistency persist when the posterior is conditioned on more than one data point per observation ($K=10$), that is, in the presence of a summary network. Further, we still see clear benefits of adding the self-consistency loss even when the likelihood is estimated by a neural likelihood approximator $q(x\mid\theta)$, trained jointly with the posterior approximator $q(\theta\mid x)$ on the same training data. However, with an estimated likelihood, posterior bias, especially bias in the posterior standard deviation, and MMD distance to the true posterior are larger than in the known likelihood case.
@@ 4.2 Forecasting air passenger traffic: an autoregressive model with predictors
Figure 3: Comparison of posterior estimates for 15 countries (ISO 3166 alpha-2 codes) among standard NPE (red circles), NPE + self-consistency loss (blue squares), and Stan (reference; gray triangles). Central 50% (thick lines) and 95% (thin lines) posterior intervals of the autoregressive component $\beta$ are shown, sorted by lower 5% quantile as per Stan (i.e., established benchmark). The self-consistency loss was evaluated on data from $M=\mathbf{8}$ countries during training, greatly enhancing ABI’s robustness in both no-misspecification scenarios and real-data evaluations.
We apply our self-consistency loss to analyze trends in European air passenger traffic data provided by Eurostat [11, 12, 13]. This case study highlights that the strong robustness gains also occur in real-world scenarios and model classes that are challenging to estimate in a simulation-based inference setting. We observe that approximators trained with the standard simulation-based loss alone yield incorrect posterior estimates for several countries. In contrast, approximators trained also with our self-consistency loss provide highly similar results to Stan as a gold-standard reference.
We retrieved time series of annual air passenger counts between 15 European countries (departures) and the USA (destination) from 2004 to 2019 and fit the following autoregressive process of order 1:
@@
where the target quantity $y_{j,t+1}$ is the difference in air passenger traffic for country $j$ between time $t+1$ and $t$. To predict $y_{j,t+1}$ we use two additional predictors: $u_{j,t}$ is the annual household debt of country $j$ at time $t$, measured in % of gross domestic product (GDP) and $w_{j,t}$ is the real GDP per capita. The parameters $\alpha_{j}$ are country-level intercepts, $\beta_{j}$ are the autoregressive coefficients, $\gamma_{j}$ are the regression coefficients of household debt and $\delta_{j}$ are the regression coefficients of GDP per capita, and $\sigma_{j}$ is the standard deviation of the noise term. This model was previously used within ABI in [24]. As commonly done for autoregressive models, we regress on time period differences to mitigate non-stationarity. This is critical for simulation-based inference because when $\beta_{j}>1$, exponential growth quickly produces unrealistic air traffic volumes. Moreover, amortizing over covariate spaces, such as varying GDP per capita between countries, can lead to model misspecification if such fluctuations are underrepresented in training. Training relies on a small simulation budget of $N=1024$, with the self-consistency loss evaluated on real data from $M\in\{4,8,15\}$ countries. Further details on training are in Appendix D.
Table 1: Posterior metrics for NPE and NPE augmented with self-consistency loss (NPE + SC) relative to Stan. For each parameter, the absolute bias in posterior means and standard deviations are reported along with the Wasserstein distance between the posteriors. The self-consistency loss was evaluated on data from $M=$ 8 countries during training. Metrics are averaged over all 15 countries.
Parameter $|\mu-\mu_{\rm Stan}|$ $|\sigma-\sigma_{\rm Stan}|$ Wasserstein distance
NPE NPE+SC NPE NPE+SC NPE NPE+SC
$\alpha$ $0.079$ 0.014 $0.033$ 0.020 $0.086$ 0.035
$\beta$ $0.153$ 0.031 $0.055$ 0.004 $0.161$ 0.054
$\gamma$ $0.087$ 0.006 $0.058$ 0.035 $0.154$ 0.068
$\delta$ $0.052$ 0.042 $0.038$ 0.031 $0.119$ 0.064
$\log(\sigma)$ $0.214$ 0.148 $0.049$ 0.011 $0.304$ 0.170
@@ Results
In Figure 3, we show exemplary results from standard NPE, our semi-supervised NPE ($M=8$), and Stan as reference. We see that standard NPE is highly inaccurate for many countries, whereas our semi-supervised approach is in strong agreement with the reference for all but one country. As shown in Table 1, adding the self-consistency loss ($M=8$) strongly improves posterior estimates for all five parameters across all metrics, on average across countries. The complete results along with standard error values for Table 1 can be found in Appendix E.
@@ 4.3 Hodgkin-Huxley model of neuron activation
To investigate the effect of the self-consistency loss on a model involving high-dimensional data, we evaluate our approach on the Hodgkin-Huxley model, which was previously used in an ABI setting by Gloeckler et al. [20]. The Hodgkin-Huxley model is a classical model in neuroscience to describe neuron activation via a set of 5 ordinary differential equations. In brief, the model has 7 parameters (electrical conductances of different ion channels, membrane capacitance and reversal potentials), and the output $y_{i}$ with observation index $i$ is a 200-dimensional time series of the membrane potential. A full definition of the model, as well as the training setup and a description of the neural architectures, are shown in Appendix F.
To facilitate network training, all parameters are transformed to follow standard normal distributions through appropriate transformations. For example, the parameter $g_{\text{Na}}$ with marginal prior $g_{\text{Na}}\sim\text{LogNormal}(\log(110),0.1²)$ is transformed via $z_{g_{\text{Na}}}=(\log(g_{\text{Na}})-\log(110))/0.1$. We denote the full set of transformed model parameters by $\theta$. Training is performed with a simulation budget of $N=32{,}768$. For each loss evaluation, the self-consistency loss is computed on a random subset of 32 samples drawn from a pool of $M=1{,}024$ unlabeled observations. These are generated by first sampling $\theta\sim\text{Normal}(0,2)$, applying the inverse transformations to recover the original parameter scale, and then simulating time series of the membrane potential as above.
Results
@@
To assess the benefits of our approach in the out-of-distribution setting, Figure 11(a) shows posterior predictive samples inferred from data simulated with $\theta\sim\text{Normal}(-2,1)$. This contrasts with training data from $\theta\sim\text{Normal}(0,1)$ and self-consistency evaluation data from $\theta\sim\text{Normal}(0,2)$. The plot shows that, when training without the self-consistency loss, the neural posterior density estimator produces samples inconsistent with the observed data, while incorporating the loss yields accurate predictions (see Appendix G for further results). To quantify this, Figure 11(b) reports mean absolute bias differences between the two estimators. The self-consistency loss consistently and strongly improves predictions across the majority of time series. Even in the worst cases, it is at least competitive with the estimator trained without the self-consistency loss.
@@ 4.4 Bayesian denoising of MNIST images
Finally, we illustrate the utility of our self-consistency loss based semi-supervised approach in the high-dimensional setting of image denoising in a set-up similar to [10]. The parameter vector $\theta\in\mathbb{R}^{784}$ is the flattened image, and the observation $x\in\mathbb{R}^{784}$ is a blurry version of the same image generated by a simulated noisy camera. We assume an implicit prior $\theta\sim p(\theta)$ defined by a generative model trained on blurred MNIST images of digit “0” [31], and an implicit likelihood $p(x|\theta)$ implemented by reapplying the same amount of blur to each $\theta$. This creates a challenging neural posterior-likelihood estimation (NPLE) problem (see also Section 2.3).
To generate our training set, we first blur all digit “0” images in the MNIST training set with a fixed Gaussian filter and train a neural network to (i) sample new blurred simulated images $\{\theta^{i}\}_{i=1}^{N}\sim p(\theta)$ and (ii) evaluate their log-probabilities. For each sample $\theta^{i}$, we then produce an observation $x^{i}\sim p(x|\theta^{i})$ by reapplying the same amount of Gaussian blur. We generate $N=12000$ pairs $(\theta^{i},x^{i})$ to train our posterior and likelihood networks. For the self-consistency loss, we use a held-out subset of 400 MNIST test set images blurred only by the likelihood model (i.e., no additional prior blur involved) during the training. This deliberate mismatch induces a prior misspecification, allowing us to evaluate the robustness gains from using self-consistency loss in an NPLE setting. More details about the training set-up and model architecture can be found in the Appendix H.
@@ Results
We perform inference on another held-out subset of MNIST test images comprising 580 images. Figure 5 (also Figure 14 in Appendix I) depicts ten randomly chosen examples alongside the posterior mean and standard deviation maps computed from 500 samples. Reconstructions using the self-consistency approach (NPLE+SC) are smoother and show more resemblance to the ground-truth image. In contrast, the means of posterior samples from NPLE are highly pixelated and blurry. Moreover, the standard deviation maps of NPLE+SC are far more coherent as elevated variations appear only along the inner and outer contours of the “0” precisely where one expects genuine edge ambiguity. In contrast, NPLE estimates exhibit scattered, patchy uncertainty across the digit and background, reflecting spurious standard deviation estimates. Figure 15 in Appendix I shows several individual posterior samples of seven randomly chosen images from the test-set for both NPLE+SC and NPLE approaches. The NPLE+SC posterior samples are smoother, less pixelated and better resemble the true image further reaffirming the advantage of including self-consistency loss.
Figure 5: Example of denoising results for MNIST images of digit “0” in the held-out test set. The first row shows ten randomly selected MNIST images ($\theta$), the second row depicts the same images after applying the Gaussian blur ($x$), third and fourth rows depict mean and SD of 500 posterior samples estimated from the corresponding blurry observations using NPLE+SC, and the fifth and sixth rows depict the mean and SD of 500 posterior samples using NPLE only. Incorporating SC loss significantly improves denoising: reconstructed means become smoother, less pixelated, and closer to the ground truth. In the standard-deviation maps, darker regions indicate higher output variability; NPLE + SC approach produces coherent maps with variability confined to the inner and outer edges.
@@ 5 Discussion
We demonstrated that Bayesian self-consistency losses significantly increase the robustness of neural amortized Bayesian inference (ABI) on out-of-simulation data. Accurate inference outside the training distribution, such as in the presence of model misspecification, has long posed a major challenge for ABI. While self-consistency was originally introduced to improve training efficiency with slow simulators [39, 26], it had not been previously explored as a remedy to simulation gaps. Existing supervised ABI approaches have been known to dramatically fail in such cases [38, 20, 25], as we also illustrated in our experiments. In contrast, when optimizing for self-consistency on unlabeled out-of-simulation data, we obtained nearly unbiased posterior estimation far beyond the training distribution. The strong robustness gains persisted even in models with several hundred parameters. Additionally, using a neural (i.e., approximate) in place of an analytic likelihood density also increased the robustness significantly. Finally, as self-consistency losses do not require data labels (i.e., true parameter values), we can use any amount of real data during training to improve the robustness of ABI.
@@
A notable limitation is that our variance-based self-consistency loss relies on fast density evaluations during training, keeping times competitive. This makes free-form methods such as flow matching [32] or score-based diffusion [41] less practical due to their need for numerical integration. As a result, efficient self-consistency losses for free-form flows, along with joint learning of posteriors and very high-dimensional likelihoods, remains an open avenue for future research.

### v5

@@ 4.1 Multivariate normal model
We first illustrate the usefulness of our proposed self-consistency loss on a controllable toy problem (Schmitt et al., 2023). The prior and likelihood are given by
$\theta\sim\text{Normal}(\mu_{\rm prior},\sigma^{2}_{\rm prior}\,I_{D}),\hskip 14.22636ptx^{(k)}\sim\text{Normal}(\theta,\sigma^{2}_{\rm lik}\,I_{D})$ (7)
The parameters $\theta\in\mathbb{R}^{D}$ are sampled from a $D$-dimensional multivariate normal distribution with mean vector $\mu_{\rm prior}$ and diagonal covariance matrix $\sigma^{2}_{\rm prior}\,I_{D}$. Here, we fix $\mu_{\rm prior}=0$ and $\sigma^{2}_{\rm prior}=1$. On this basis, $K$ independent, synthetic data points $x^{(k)}\in\mathbb{R}^{D}$ are sampled from a $D$-dimensional multivariate normal distribution with mean vector $\theta$ and diagonal covariance matrix $\sigma^{2}_{\rm lik}\,I_{D}$. We set $\sigma^{2}_{\rm lik}=K$ such that the total information in $x$ remains constant across values of $K$, which simplifies comparisons across observations of varying numbers of data points. Refer to Appendix C for more details on training and neural architectures.
@@ Results
Figure 2: Posterior distance quantified by maximum mean discrepancy (MMD) to the analytic posterior for variations of the default configuration. Errorbars show $\pm 1$ SDs over 10 model refits.
In Figure 1, we depict the results obtained from (a) standard NPE (trained on the simulation-based loss only), (b) our semi-supervised NPE + SC (with self-consistency loss on known likelihood), and (c) the gold-standard (analytic) reference. We see that standard NPE already completely fails for $x_{\rm obs}\sim N(\mu_{\rm obs}=2,0.01I_{D})$, and subsequently also for any larger values $\mu_{\rm obs}>2$. In contrast, our semi-supervised approach achieves almost perfect posterior estimation. This holds true even in cases where $x_{\rm obs}$ is multiple standard deviations away from all the training data, that is, from both the labeled dataset $\{(\theta_{n},x_{n})\}_{n=1}^{N}$ and the unlabeled dataset $\{{x^{*}_{m}}\}_{m=1}^{M}$. These results indicate that the self-consistency criterion can provide strong robustness gains even far outside the typical space of training data.
In Figure 2, we report the maximum mean discrepancy (MMD) between the approximate and true posterior the factors parameter dimensionality and number of unlabeled observations. When varying the parameter dimensionality (Figure 2(a)), including the SC loss yields nearly perfect posterior approximation up to 10 dimensions, even with extreme deviations from the initial training data. It also significantly improves accuracy in the 100 parameter scenario. The dataset size factor (Figure 2(b)) shows robust gains, with clear improvements over NPE even when using as few as four unlabeled observations (versus 1024 labeled ones). In Figure 6 (Appendix D), we report posterior mean and standard deviation bias as well as maximum mean discrepancy for all the above factors. Varying the mean $\mu_{\rm obs}$ of the new observations shows that, as long as the data used for evaluating the SC loss is not identical to the training data (i.e., as long as $\mu_{\rm obs}\neq 0$), including the SC loss component enables accurate posterior approximation far outside the typical space of the training data.
In Figure 7 in Appendix D, we see that the benefits of self-consistency persist when the posterior is conditioned on more than one data point per observation ($K=10$), that is, in the presence of a summary network. Further, we still see clear benefits of adding the self-consistency loss even when the likelihood is estimated by a neural likelihood approximator $q(x\mid\theta)$, trained jointly with the posterior approximator $q(\theta\mid x)$ on the same training data. However, with an estimated likelihood, posterior bias, especially bias in the posterior standard deviation, and MMD distance to the true posterior are larger than in the known likelihood case. We also compared our semi-supervised approach with other robustness methods - NPE-DANN, NPE-MMD, and NNPE. The results described in Appendix D.1 demonstrate the benefits of our approach over other methods.
@@ 4.2 Forecasting air passenger traffic: an autoregressive model with predictors
We apply our semi-supervised loss to analyze trends in European air passenger traffic data provided by Eurostat (2022b, a, c). This case study highlights that the strong robustness gains also occur in real-world scenarios and model classes that are challenging to estimate in a simulation-based inference setting. We retrieved time series of annual air passenger counts between 15 European countries (departures) and the USA (destination) from 2004 to 2019 and fit the following autoregressive process of order 1:
@@
where the target quantity $y_{j,t+1}$ is the difference in air passenger traffic for country $j$ between time $t+1$ and $t$. To predict $y_{j,t+1}$ we use two additional predictors: $u_{j,t}$ is the annual household debt of country $j$ at time $t$, measured in % of gross domestic product (GDP) and $w_{j,t}$ is the real GDP per capita. Training relies on a small simulation budget of $N=1024$, with the self-consistency loss evaluated on real data from $M\in\{4,8,15\}$ countries. Further details on model and training are provided in Appendix E.
@@ Results
Figure 3: Comparison of posterior estimates for 15 countries (ISO 3166 alpha-2 codes) among standard NPE (red), NPE + SC (blue), and Stan (reference in gray; Carpenter et al., 2017). Central 50% (thick lines) and 95% (thin lines) posterior intervals of the autoregressive component $\beta$ are shown, sorted by lower 5% quantile as per Stan (i.e., established benchmark). The SC loss was evaluated on data from $M=\mathbf{8}$ countries during training, greatly enhancing ABI’s robustness in both no-misspecification scenarios and real-data evaluations.
In Figure 3, we show exemplary results from standard NPE, our semi-supervised NPE + SC ($M=8$), and Stan (Carpenter et al., 2017) as reference. We see that NPE is highly inaccurate for many countries, whereas NPE + SC is in strong agreement with the reference for all but one country. As shown in Table 1, adding the self-consistency loss ($M=8$) strongly improves posterior estimates for all five parameters across all metrics, on average across countries. The complete results including comparisons of NPE+SC with online trained NPE, VAE, and Schmitt et al. (2024) can be found in Appendix F and F.1.
@@ 4.3 Hodgkin-Huxley model of neuron activation
To investigate the effect of the self-consistency loss on a model involving high-dimensional data, we evaluate our approach on the Hodgkin-Huxley model, which was previously used in an ABI setting by Gloeckler et al. (2023). The Hodgkin-Huxley model is a classical model in neuroscience to describe neuron activation via a set of 5 ordinary differential equations. In brief, the model has 7 parameters (electrical conductances of different ion channels, membrane capacitance and reversal potentials), and the output $y_{i}$ with observation index $i$ is a 200-dimensional time series of the membrane potential. A full definition of the model, as well as the training setup and a description of the neural architectures, are provided in Appendix G.
To facilitate network training, all parameters are transformed to follow standard normal distributions through appropriate transformations. For example, the parameter $g_{\text{Na}}$ with marginal prior $g_{\text{Na}}\sim\text{LogNormal}(\log(110),0.1²)$ is transformed via $z_{g_{\text{Na}}}=(\log(g_{\text{Na}})-\log(110))/0.1$. We denote the full set of transformed model parameters by $\theta$. Training is performed with a simulation budget of N=32,768N=32{,}768. For each loss evaluation, the self-consistency loss is computed on a random subset of 32 samples drawn from a pool of M=1,024M=1{,}024 unlabeled observations. These are generated by first sampling $\theta\sim\text{Normal}(0,2)$, applying the inverse transformations to recover the original parameter scale, and then simulating time series of the membrane potential as above.
Results
To assess the benefits of our approach in the out-of-distribution setting, Figure 4(a) shows posterior predictive samples inferred from data simulated with $\theta\sim\text{Normal}(-2,1)$. This contrasts with training data from $\theta\sim\text{Normal}(0,1)$ and self-consistency evaluation data from $\theta\sim\text{Normal}(0,2)$. When training without the self-consistency loss, the neural posterior density estimator produces samples inconsistent with the observed data, while incorporating the loss yields accurate predictions (see Appendix H for further results). To quantify this, Figure 4(b) reports mean absolute bias differences between the two estimators. The self-consistency loss consistently and strongly improves predictions across the majority of time series. Even in the worst cases, it is at least competitive with the estimator trained without the self-consistency loss.
@@ 4.4 Bayesian denoising of MNIST images
Figure 5: Example of denoising results for MNIST images of digit “0” in the held-out test set. The first row shows ten randomly selected MNIST images ($\theta$), the second row depicts images after applying the Gaussian blur ($x$), third and fourth rows depict mean and standard deviation (SD) of 500 posterior samples estimated from the corresponding blurry observations using NPLE+SC, and the fifth and sixth rows depict the mean and SD of 500 posterior samples using NPLE only. Incorporating SC loss significantly improves denoising.
Finally, we demonstrate the utility of our self-consistency loss based approach in a high-dimensional (for ABI) image denoising task, following a set-up similar to Elsemüller et al. (2025). The parameter vector $\theta\in\mathbb{R}^{784}$ is the flattened image, and the observation $x\in\mathbb{R}^{784}$ is a blurred version generated by a simulated noisy camera. We assume an implicit prior $\theta\sim p(\theta)$ defined by a generative model trained on blurred MNIST images of the digit “0” (LeCun et al., 1998), and an implicit likelihood $p(x|\theta)$ implemented by reapplying the same blur to each $\theta$. This creates a challenging neural posterior-likelihood estimation (NPLE) problem.
To construct the training set, we blur all digit “0” images in the MNIST training set with a fixed Gaussian filter and train a neural network to (i) sample blurred simulated images $\{\theta^{i}\}_{i=1}^{N}\sim p(\theta)$ and (ii) evaluate their log-probabilities. For each sample $\theta^{i}$, we then generate an observation $x^{i}\sim p(x|\theta^{i})$ by reapplying the same Gaussian blur, yielding $N=12000$ pairs $(\theta^{i},x^{i})$ to train posterior and likelihood networks. For the SC loss, we use 400 held-out MNIST test images blurred only by the likelihood model, (i.e., no additional prior blur), inducing deliberate prior misspecification and enabling us to asses the robustness of SC in NPLE. More details about the training set-up and model architecture can be found in the Appendix I.
@@ Results
We perform inference on a separate subset of 580 MNIST test images. Figure 5 and Figure 16 in Appendix J show randomly chosen examples alongside the posterior mean and standard deviation maps computed from 500 posterior samples. Reconstructions using the self-consistency approach (NPLE+SC) are smoother and more faithful to the ground-truth, whereas NPLE reconstructions are pixelated and blurry. Moreover, NPLE+SC yields coherent uncertainty maps, with elevated variance localized along digit contours where edge ambiguity is expected. In contrast, NPLE estimates exhibit scattered, patchy uncertainty across the digit and background. Additional examples are reported in Appendix J.
@@ 5 Discussion
We demonstrated that Bayesian self-consistency losses significantly increase the robustness of neural amortized Bayesian inference (ABI) on out-of-simulation data. Accurate inference outside the training distribution, such as in the presence of model misspecification, has long posed a major challenge for ABI. While self-consistency was originally introduced to improve training efficiency with slow simulators (Schmitt et al., 2024; Ivanova et al., 2024), it had not been previously explored as a remedy to simulation gaps. Existing supervised ABI approaches have been known to dramatically fail in such cases (Schmitt et al., 2023; Gloeckler et al., 2023; Huang et al., 2023), as we also illustrated in our experiments. In contrast, when optimizing for self-consistency on unlabeled out-of-simulation data, we obtained nearly unbiased posterior estimation far beyond the training distribution. The strong robustness gains persisted even in models with several hundred parameters. However, using an approximate (neural) likelihood has not yet matched the robustness of the known-likelihood case. Finally, as self-consistency losses do not require data labels (i.e., true parameter values), we can use any amount of real data during training to improve the robustness of ABI.
@@
Our variance-based self‐consistency loss relies on fast density evaluations during training, keeping times competitive. This makes free‐form methods such as flow matching (Lipman et al., 2023) or score-based diffusion (Song et al., 2020) less practical due to their need for numerical integration. As a result, efficient self‐consistency losses for free‐form flows, along with joint learning of posteriors and very high-dimensional likelihoods, remains an open avenue for future research.
Our semi-supervised approach can also be combined with post-hoc correction methods such as Pareto-smoothed importance sampling (Vehtari et al., 2024). When applied to NPE alone, importance sampling often fails because the learned posterior can deviate substantially from the analytic target (Li et al., 2024). By contrast, our approach produces posterior approximations that are much closer to the analytic posterior, making subsequent importance sampling more effective. This combination therefore has the potential to substantially expand the range of scenarios in which accurate estimation of the analytic posterior is feasible.